Astronomy Report – December 30th

Name: Cesare Guariniello

Crew: 218

Date: 30-12-2019


Submitted an observation of M42 (Orion Nebula, 10 exposures of 1 second each in B, V, R) and M31 (Andromeda Galaxy, 10 exposures of 20 seconds each in B, V, R)

Images submitted with this report: M42 29122019 and M31 29122019

Problems encountered: working on finding the best exposure time. Red flat correction not applicable, makes the image worse.

GreenHab Report – December 30th

Crew 218 GreenHab Report 30-DEC-19

Crew GreenHab Officer: Dr. Jonathan R. Buzan

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 23.9°C; 20%


Floor Unit: 15°C

Electronic: 18.7°C

humidity 21%


Floor Unit: 16°C

Electronic: 20.9°C

humidity 20%


Floor Unit: 15°C

Electronic: 22.2°C

humidity 19%

Max: 25.5; 21%

Min: 22.3°C; 19%

Hours of supplemental light: Light system 6:00PM-11:00AM.

Daily water usage for crops: 6.5Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – ~93 Gal.

Time(s) of watering for crops:

2:20PM watered hanging plants by heater with miracle grow (2.2Gal.).

6:00PM watered with miracle grow (4.3Gal.).


Change to crops: None.


1st sprouts: N/A

—2:20PM Cucumber was dry and wilting. The plant has flowered. Does flowering consume more water? Watered the cucumber and hanging plants by heater with MiracleGrow.

—One Mystery Plant has flowered.

—Acorn Pumpkin needs to be transplanted to larger pot.


Cilantro: 3.95g

Basil: 2.00g

Carrot Tops: 3.25g

Crew 218 Journalist Report 30Dec2019

Journalist Report
Ben Durkee, Crew 218 Journalist

Sol 08

The snowy hill east of the Habitat delayed the sunrise just long enough for us to awaken to a sky of vibrant pink and orange. Nothing clears the morning haze from your eyes quite like gawking at the natural beauty of a Martian sunrise. We were fortunate to have such a gentle wake-up call, because the rest of the day was on a tight itinerary.

After an efficient breakfast of astronaut nutrition paste (dehydrated fruit smoothie) we promptly began preparing Shefali and LuzMa for their morning EVA. After an egress that went infinitely smoother than yesterday’s, they set off into the icy brink. The two ladies had the honor of taking our rovers "T" and "Custy" (formerly "Spirit" and "Curiosity," but some of the letters rubbed off) on their first voyage in a week.

The rovers persevered through the quagmire of snow, ice, and mud to deliver the duo to their destination unscathed. The two performed some meteorological surveys that are far beyond my pay-grade and returned to the Hab far ahead of schedule. In an effort to kill two birds with one Martian stone, LuzMa also performed her extravehicular engineer duties before they both entered the airlock for re-compression. Everything went off without a hitch, and they were inside and free of their cumbersome oxygen-backpacks with time to spare.

More time to dedicate to the most important task of the day. See, the past few sols have been consistently below freezing, and our water pipe has been an ice pipe for long enough that we are in full survival mode. The first thing to go when the water gets tight is dish duty, and by today we had a tower rivaling the wonders of the world constructed solely of dirty dishes. Our water reserve was so sparse this afternoon that we had to resort to creating a fire brigade. We spent the period between EVAs today assembly-lining water from the static tank all the way up to the accessible loft tank with our few clean kitchen pots. Our efficiency would have made Henry Ford proud – probably because we had pretty compelling motivation.

Once our liquid life reached an acceptable quantity, there was just enough time for a quick lunch before Cesare, Pat, and I had to prepare for our EVA. We raced to see who could be properly suited up first, and as expected our commander with multiple notches in his Mars belt pulled a clean victory. This time I abundantly confirmed that I was wearing the right footwear. I’m sure I’ll make many more mistakes this mission, but I will definitely not be making that one again.

While we were acting as a human aqueduct, our rovers were replenishing their batteries back to full. For good reason – we intended to use it all. We depressurized and embarked on our longest EVA yet. Our destination: as far north as we could go.

There is a clear division drawn in mud on my flight suit. A contrast between my left side: cozy within the roll cage of trusty ol’ "Custy," and my right side: exposed to the elements as I held out the antenna for some data collection on our northward exodus. We adventured far into the Martian horizon on a journey that felt like an eternity for the muscles in my right arm. We exited the zone of radio contact with the Hab, and then we pushed on even farther. Eventually Cesare’s rover read 60% battery remaining, our threshold for having to cease our migration. We pulled over right there, and the spot proved to be ideal.

It had a flat region – perfect for Pat’s seismic equipment – and a phenomenal view of the nearby mountain – a playground for me and my camera. We spent equal time setting up the equipment as we did gathering data. With our remaining time, we explored the surrounding terrain and then gave our rovers a good 20-point turn to head back home.

Naturally the journey somehow managed to be uphill both ways, but the rovers took it like champs and we were back in no time. By the time we got ourselves and our equipment back in the airlock, it was 4:00 PM sharp. A perfect three hour EVA. We raced out of our suits again, but this time it was fueled by the necessity to empty our bladders. We all parsed our data from the field and then put on some SpongeBob to turn our brains off and recover from the day. Much needed.

Through the yellow rectangle man’s uncanny ability to evaporate time, it was suddenly time for dinner and reports. As Pat slaves away in the kitchen, I’m hunched over my keyboard relaying the day’s events to the best of my ability. It took a week, but I think we’re getting the hang of this whole "Martian Living" thing. Now we wrap up today so we can do it again tomorrow, but even better.

Another sol, another dollar.

Crew 218 EVA Report 30Dec2019

EVA #: 4

Author: Shefali Rana

Purpose of EVA: weather observation project; EVA and EMU evaluation project

Start time: 10:10

End time: 11:22

Narrative: Luz Ma and Shefali took Curiosity and Spirit up to the Galileo Rd intersection and took observation readings for weather project. The EVA duration was about one hour. After being back in the habitat, Luz Ma and Shefali filled questionnaires for the EVA and EMU evaluation projects.

Destination: Galileo Rd

Coordinates (use UTM NAD27 CONUS): 12S 519050, 4252000

EVA Participants: LuzMa Agudelo, ENG; Shefali Rana, HSO

Road(s) and routes per MDRS Map: Cow Dung Road

Mode of travel: Driving, then work in the field and walking

EVA #: 5

Author: Cesare Guariniello

Purpose of EVA: crew decision-making evaluation and radio measurements (during transfers); seismometric measurements to evaluate subsurface layers (once at destination); EVA and EMU evaluation (once back in habitat)

Start time: 13:15

End time: 15:59

Narrative: After loading on the rovers the necessary equipment for the research projects, we began driving along Cow Dung Road. The road is well marked, with streams of icy water melting down. Snow is completely melted in certain point and about one inch deep on others. The ride had been very smooth and safe, and the rovers made it to the intended destinations with good residual charge. Ben took radio measurement along the road, as a passenger on Curiosity rover. Once at destination, we began to slowly prepare Pat’s experiment. This required us to extend a measuring tape for about 30 meters and, after having removed some snow, place 24 sensors one meter apart from each other. The seismometric sensors, connected to a laptop, record the propagation of seismic waves generated by a hammer hitting a metal plate that we laid on the ground at various locations along the sensor network. The preparation took about 25 minutes and gave us a good idea of the challenges faced by astronauts working in EVA. After the experiment, we walked the first few feet of Brahe Hwy before heading back. On our way back, Ben took more radio measurements south of the intersection with Galileo Rd and close to the habitat.

Destination: Brahe Hwy at Glistening Sea

Coordinates (use UTM NAD27 CONUS): 12S 517750, 4254700

EVA Participants: Cesare Guariniello, CMD; Pat Pesa, GEO; Ben Durkee, JOU

Road(s) and routes per MDRS Map: Cow Dung Road to Brahe Hwy intersection (Glistening Sea and White Moon area)

Mode of travel: Driving and walking

Crew 218 Science Report 30Dec2019

[title Science Report – December 30th]

[category science-report]

Crew 218 Science Report 30-DEC-19
Crew Science Officer: Dr. Jonathan R. Buzan

1. Decision Making in support of autonomy for crew EVAs:

No EVA: nothing to report.

2. Mars surface weather:

Measurements were collected in yesterday’s and today’s EVA.

3. Subsurface structure on Mars:

The seismic reflection mapping instrument was deployed in today’s EVA.

4. Detecting radio signal strength:

Field measurements were collected in yesterday’s and today’s EVAs.

5. EVA workload analysis:

Survey data was collected for yesterday’s and today’s EVAs.

6. EMU (Extravehicular Mobility Unit) ergonomic assessment:

Survey data was collected for yesterday’s and today’s EVAs.

7. Environmental Stresses over MDRS habitat and Crew Members and projection over Martian Terrain:

Nothing to report.

8. Messier and other space objects for outreach:

Two astronomy photos were taken.

9. Reliability and maintenance:

No EVA: nothing to report.

10. Medical readings in preparation for future crew-wide project:

The Zephyr bio metric monitor collected data for 20 hours. Battery was changed and used for EVA today.

11. Collection of clay, shale, and hematite samples:

3 rock samples were collected.

12. Media and outreach:

Nothing to report.

Glassware check out: None

A reminder to all crewmembers: There’s a $300 fine for using any glassware material without MDRS Mission support permission.

Crew 218 Operations Report 30Dec2019

[title Operations Report – December 30th]

[category operations-report]

Crew 218 Operations Report 30-Dec-19

SOL: 08

Name of person filing report: Luz Maria Agudelo Urrego

Non-nominal systems: NA

Generator: run

Hours run: 14

From what time last night: 1730

To what time this morning: 0730

List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before the generator is run at night: 75%

Diesel Reading – 76%

Station Propane Reading – 69%

Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 42

Water Meter (units): 0147145.7

Water (static tank) (gal): 345

Static to Loft Pump used – No

Water in Green Hab (gal): 93

Water in Science Dome (gal): 0

Toilet tank emptied: No

Deimos rover used: Still in the workshop


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: Assigned to director


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Nominal

Hours: 120.3

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging): 30%

Currently charging: Yes

Opportunity rover used: Still in the workshop


Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: Nominal

Hours: 125.5

Beginning charge: 100%

Ending charge: (On return from EVA, before charging): 29%

Currently charging: Yes

Notes on rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use: N/A

Oil Added? No

# Hours the ATVs were used today: 0

Notes on ATVs: N/A

HabCar used, and why, where? No

CrewCar used, and why, where? No

General notes and comments: N/A

Summary of the internet: Nominal

Summary of suits and radios: Nominal

Summary of Hab operations: The pipe from the static tank through the loft tank is still frozen. The crew filled out the Lott Tank by carrying water buckets from the Static Tank.

Summary of Science Dome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: Nominal

Summary of health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: NA

Commander Report – December 30th

Crew 218 Commander Report – 30-12-2019

Sol 8 – Out on white Mars

Today we had a long and productive day. Our dishes piling up because of frozen pipes, we opted for a quick smoothie for breakfast before delving into EVA activity. Luz Ma and Shefali had our first all-women EVA, destination Galileo Road, and they surveyed the situation of the roads for us. Luz Ma took weather observations while Shefali could participate in person in her project of evaluation of EVA and EMU suits. Once they came back, we completed a team activity to use an auxiliary pump, a hose, and various pots to transfer water from our large static tank to the loft tank, that is the source of water for our kitchen, washroom, and toilet. The second EVA, with me, Pat and Ben, took a longer route and brought us to the intersection with Brahe Hwy. All three of us collected valuable data for our projects. When we re-entered, we were welcome by fresh, warm rosemary bread which I had started before EVA egress. Reinforced by the various adversities we faced together, we are on a steep learning curve to become better Martians inside and outside the habitat!

Sol Summary – December 30th

Sol: 08

Summary Title: Martian Explorers

Author’s name: Pat Pesa

Mission Status: Crew is feeling accomplished!

Sol Activity Summary: Crew 218 had many sources of excitement today. We had two great EVAs that allowed for significant progress in progress in the Climate Monitoring, Detecting Radio Signals, EVA suit ergonomics, as well as starting the Geologist’s project of analyzing Sub-Surface Ground Structure.

Outside of our EVAs we also "solved" our water troubles (at least delayed them a day or two) by manually filling our Loft tank via human chain and buckets of water from the static tank (fireman’s style).

Look Ahead Plan: More EVAs! Also tomorrow will be New Years Eve, so we have some fun activities planned for the evening.

Anomalies in work: None

Weather: Partly Cloudy, around 32 degrees

Crew physical status: Healthy

EVA: EVA #4 this morning went to Galileo Road to take Climate data. EVA #5 this afternoon went nearly to the crossing of Brahe Highway (we had to stop slightly before due to Rover Battery levels), taking continuous Radio Signal Measurements, and a Seismic Survey for the Ground Structure Project.

Reports to be filed: sol summary, commander report, operations report, greenhab report, journalist report, EVA report, EVA request, science report

Support Requested: none

Copyright © The Mars Society. All rights reserved. | Main Site